# Thresholded Lasso for high dimensional variable selection

### v2

@@ Appendix D Proof of the MSE lower bound
$\left\lVert\hat{\beta}-\beta\right\rVert_{2}$$\leq$ $\left\lVert\hat{\beta}-\beta_{T_{0}}\right\rVert_{2}+\left\lVert\beta-\beta_{T_{0}}\right\rVert_{2}\leq\left\lVert h\right\rVert_{2}+\lambda\sigma\sqrt{s_{0}}$
@@
Appendix I Nearly ideal model selections under the UUP
@@
Remark I.2.
It is clear that $\delta_{2s}<1$ implies that the sparse eigenvalues condition (6) and (5) hold. Moreover, Assumption I.1 implies that $\textnormal{{RE}}(s_{0},k_{0},X)$ as in (3) hold for $s_{0}\leq s$ with
@@
Theorem I.3.
(Variable selection under UUP) Choose $\tau,a>0$ and set $\lambda_{n}=\lambda_{p,\tau}\sigma$, where $\lambda_{p,\tau}:=(\sqrt{1+a}+\tau^{-1})\sqrt{2\log p/n}$, in (153). Suppose $\beta$ is $s$-sparse with $\delta_{2s}+\theta_{s,2s}<1-\tau$. Let threshold $t_{0}$ be chosen from the range $(C_{1}\lambda_{p,\tau}\sigma,C_{4}\lambda_{p,\tau}\sigma]$ for some constants $C_{1},C_{4}$ to be defined. Then the Gauss-Dantzig selector $\hat{\beta}$ selects a model $I:=\mathop{\text{\rm supp}\kern 0.86108pt}(\hat{\beta})$ such that we have
@@
Our analysis for Theorem I.3 builds upon Proposition I.4 [9], which shows the Dantzig selector achieves the oracle inequality as stated in (11) under Assumption I.1. We note that, in Assumption I.1, the sparsity level is fixed at $s$ rather than $s_{0}$. Hence it is stronger than the conditions we impose in Theorem 2.1 for the Thresholded Lasso estimator. We now show the oracle inequalities for the Dantzig selector. We then show in the supplementary Lemma J.1 that thresholding at the level of $\sigma\lambda$ as elaborated in Step 1 in the Gauss-Dantzig Procedure selects a set $I$ of at most $2s_{0}$ variables, among which at most $s_{0}$ are from the complement of the support set $S$ as required in (155).
Proposition I.4.
@@
Appendix J Proof of Theorem I.3
Now similar to Lemma 2.3, Lemma J.1 bounds the size of $I$, as well as the bias that we introduce to model $I$ by thresholding. Theorem I.3 is an immediate corollary of Lemmas 2.5 and J.1. The proof follows from Lemma J.1 and Proposition I.4. We include its proof in Section J.1 for self-containment.
Lemma J.1.
@@
Proof of Theorem I.3. It holds by definition of $S_{{\mathcal{D}}}$ that $I\cap S_{{\mathcal{D}}}=\emptyset$. It is clear by Lemma J.1 that $|{S_{{\mathcal{D}}}}|<s$ and $|I|\leq 2s_{0}$ and $|I\cup S_{{\mathcal{D}}}|\leq|I\cup S|\leq s+s_{0}\leq 2s$; Thus for $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$ and $\lambda=\sqrt{2\log p/n}$, we have by Lemma 2.7
$\left\lVert\hat{\beta}_{I}-\beta\right\rVert_{2}^{2}\leq\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}\left(1+\frac{2\theta^{2}_{s,2s_{0}}}{\Lambda_{\min}^{2}(2s_{0})}\right)+\frac{4s_{0}}{\Lambda_{\min}^{2}(2s_{0})}\lambda_{\sigma,a,p}^{2}$
@@
The proof of Proposition I.4 [9] yields the following on ${\mathcal{T}}_{a}$,
@@
The rest of the proof follows that of Lemma 2.3 and hence omitted.
@@
Figure 5: Illustrative example: i.i.d. Gaussian ensemble; $p=256$, $n=72$, $s=8$, and $\sigma=\sqrt{s}/3$. (a) compare with the Lasso estimator $\tilde{\beta}$ which minimizes $\ell_{2}$ loss. Here $\tilde{\beta}$ has only 3 FPs, but $\rho^{2}$ is large with a value of 64.7364.7364.73. (b) Compare with the $\beta_{\text{\rm init}}$ obtained using $\lambda_{n}$. The dotted lines show the thresholding level $t_{0}$. The $\beta_{\text{\rm init}}$ has 15 FPs, all of which were cut after the 2nd step; resulting $\rho^{2}=12.73$. After refitting with OLS in the 3rd step, for the $\hat{\beta}$, $\rho^{2}$ is further reduced to 0.510.510.51.
Appendix K Numerical experiments
@@
The experiment set up here is the same as the one in Section  5.1. The main difference is that all non-zero entries in $\beta$ have large magnitudes around 1, in particular, they follow this distribution $\beta_{i}=\mu_{i}(1+|g_{i}|),$ where $\mu_{i}=\pm 1$ with probability 1/2 and $g_{i}\sim N(0,1)$. We use $\lambda_{n}=0.69\lambda\sigma$ throughout the experiments in this section to select a $\beta_{\text{\rm init}}$ as the initial estimator. We then threshold the $\beta_{\text{\rm init}}$ using a value $t_{0}$ typically chosen between $0.5\lambda\sigma$ and $\lambda\sigma$. See each experiment for the actual value used. Given that columns of $X$ being normalized to have $\ell_{2}$ norm $\sqrt{n}$, for each input $\beta$, we compute its SNR as follows: $SNR:=\left\lVert\beta\right\rVert_{2}^{2}/\sigma^{2}.$ To evaluate $\hat{\beta}$, we use metrics defined in Table 4; we also compute the ratio between squared $\ell_{2}$ error and the ideal mean squared error, known as the $\rho^{2}$; see Section K.2 for details.
K.1 Type I/II errors
We now evaluate the Thresholded Lasso estimator by comparing Type I/II errors under different values of $t_{0}$ and SNR. We consider Gaussian random matrices for the design $X$ with both diagonal and Toeplitz covariance. We refer to the former as i.i.d. Gaussian ensemble and the latter as Toeplitz ensemble. In the Toeplitz case, the covariance is given by $T(\gamma)_{i,j}=\gamma^{|i-j|}$ where $0<\gamma<1$ We run under two noise levels: $\sigma=\sqrt{s}/3$ and $\sigma=\sqrt{s}$. For each $\sigma$, we vary the threshold $t_{0}$ from $0.01\lambda\sigma$ to $1.5\lambda\sigma$. For each $\sigma$ and $t_{0}$ combination, we run the experiment as described in Section 5.1 200200200 times with a new $\beta$ and $\epsilon$ generated each time and we count the number of Type I and II errors in $\hat{\beta}$. We compute the average at the end of 200 runs, which will correspond to one data point on the curves in Figure 6 (a) and (b).
@@
False Positive Rate (FPR) $FPR=FP/(FP+TN)=FP/(p-s)$
@@
Figure 6: $p=256$ $s=8$. (a) (b) Type I/II errors for i.i.d. Gaussian and Toeplitz ensembles. Each vertical bar represents ±1$\pm$ 1± 1 std. The unit of $x$-axis is in $\lambda\sigma$. For both types of design matrices, FPs decrease and FNs increase as the threshold increases. For Toeplitz ensembles, in (c) with fixed correlation $\gamma$, FNs decrease with more samples, and in (d) with fixed sample size, FNs decrease as the correlation $\gamma$ decreases. (e) (f) Histograms of $\rho^{2}$ under i.i.d Gaussian ensembles from 500 runs.
We first run the above experiment using i.i.d. Gaussian ensemble under the following thresholds: $t_{0}=\lambda\sigma$ for $\sigma=\sqrt{s}/3$, and $t_{0}=0.36\lambda\sigma$ for $\sigma=\sqrt{s}$. These are chosen based on the desire to have low errors of both types (as shown in Figure 6 (a)). Naturally, for low SNR cases, small $t_{0}$ will reduce Type II errors. In practice, we suggest using cross-validations to choose the exact constants in front of $\lambda\sigma$; See, for example, a subsequent work [50] for details. We plot the histograms of $\rho^{2}$ in Figure 6 (e) and (f). In (e), the mean and median are 1.451.451.45 and 1.011.011.01 for the Thresholded Lasso, and 46.9746.9746.97 and 41.1241.1241.12 for the Lasso. In (f), the corresponding values are 7.267.267.26 and 6.606.606.60 for the Thresholded Lasso and 10.5010.5010.50 and 10.0110.0110.01 for the Lasso. With high SNR, the Thresholded Lasso performs extremely well; with low SNR, the improvement of the Thresholded Lasso over the ordinary Lasso is less prominent; this is in close correspondence with the Gauss-Dantzig selector’s behavior as shown by [9].
Next we run the above experiment under different sparsity values of $s$. We again use i.i.d. Gaussian ensemble with $p=2000$, $n=400$, and $\sigma=\sqrt{s}/3$. The threshold is set at $t_{0}=\lambda\sigma$. The SNR for different $s$ is fixed at around 32.3632.3632.36. Table 5 shows the mean of the $\rho^{2}$ for the Lasso and the Thresholded Lasso estimators. The Thresholded Lasso performs consistently better than the ordinary Lasso until about $s=80$, after which both break down. For the Lasso, we always choose from the full regularization path the optimal $\tilde{\beta}$ that has the minimum $\ell_{2}$ loss.
@@
K.3 Linear Sparsity
[…]
We experiment with $\sigma=1$ and $\sigma=\sqrt{s}/3$. The results are shown in Figure 7. We observe that under both noise levels, the Thresholded Lasso estimator requires much fewer samples than the ordinary lasso in order to conduct exact recovery of the sparsity pattern of the true linear model when all non-zero components are sufficiently large. When $\sigma$ is fixed as $s$ increases, the SNR is increasing; the experimental results illustrate the behavior of sparse recovery when it is close to the noiseless setting. Given the same sparsity, more samples are required for the low SNR case to reach the same level of success rate. Similar behavior was also observed for Toeplitz and Bernoulli ensembles with i.i.d. ±1$\pm$ 1± 1 entries.
@@
Figure 7: (a) (b) Compare the probability of success for $p=256$ and $p=512$ under two noise levels. The Thresholded Lasso estimator requires much fewer samples than the ordinary Lasso. (c) (d) (e) show the probability of success of the Thresholded Lasso under different levels of sparsity and noise levels when $n$ increases for $p=512$ and 102410241024. (f) The number of samples $n$ increases almost linearly with $s$ for p = 1024. More samples are required to achieve the same level of success when $\sigma=\sqrt{s}/3$ due to the relatively low SNR.
K.4 ROC comparison
We now compare the performance of the Thresholded Lasso estimator with the Lasso and the Adaptive Lasso by examining their ROC curves. Our parameters are $p=512$, $n=330$, $s=64$ and we run under two cases: $\sigma=\sqrt{s}/3$ and $\sigma=\sqrt{s}$. In the Thresholded Lasso, we vary the threshold level from $0.01\lambda\sigma$ to $1.5\lambda\sigma$. For each threshold, we run the experiment described in Section K.1 with an i.i.d. Gaussian ensemble $X$ being fixed while repeating Steps 232-32 - 3 100 times. After each run, we compute the FPR and TPR of the $\hat{\beta}$, and compute their averages after 100 runs as the FPR and TPR for this threshold. For the Lasso, we compute the FPR and TPR for each output vector along its entire regularization path. For the Adaptive Lasso, we use the optimal output $\tilde{\beta}$ in terms of $\ell_{2}$ loss from the initial Lasso penalization path as the input to its second step, that is, we set $\beta_{\text{\rm init}}:=\tilde{\beta}$ and use $w_{j}=1/\beta_{\text{\rm init},j}$ to compute the weights for penalizing those non-zero components in $\beta_{\text{\rm init}}$ in the second step, while all zero components of $\beta_{\text{\rm init}}$ are now removed. We then compute the FPR and TPR for each vector that we obtain from the second step’s LARS output. We implement the algorithms as given in  [52], the details of which are omitted here as its implementation has become standard. The ROC curves are plotted in Figure 8. The Thresholded Lasso performs better than both the ordinary Lasso and the Adaptive Lasso; its advantage is more apparent when the SNR is high.

### v3

@@ Appendix A Proof of the MSE lower bound
$\left\lVert\hat{\beta}-\beta\right\rVert_{2}$ $\leq$ $\left\lVert h\right\rVert_{2}+\lambda\sigma\sqrt{s_{0}}$
@@
Appendix F Proof of Lemma D.2
Thus it follows from the proof Lemma 3.1 that
$\left\lVert h_{T_{01}}\right\rVert_{2}$ $\leq$ $\left\lVert Xh_{T_{01}}\right\rVert_{2}/\sqrt{n\Lambda_{\min}(2s_{0})}$ (152)
where we replace $\frac{\theta_{s_{0},2s_{0}}}{\sqrt{\Lambda_{\min}(2s_{0})}}$ with $\sqrt{\Lambda_{\max}(s_{0})}$. The lemma thus holds in view of (38) and (152). ∎
Appendix G Nearly ideal model selections under the UUP
@@
Remark G.2.
It is clear that $\delta_{2s}<1$ implies that the sparse eigenvalues condition (6) and (5) hold. Moreover, Assumption G.1 implies that $\textnormal{{RE}}(s_{0},k_{0},X)$ as in (3) hold for $s_{0}\leq s$ with
@@
Theorem G.3.
[…]
@@
Our analysis for Theorem G.3 builds upon Proposition G.4 [9], which shows the Dantzig selector achieves the oracle inequality as stated in (11) under Assumption G.1. We note that, in Assumption G.1, the sparsity level is fixed at $s$ rather than $s_{0}$. Hence it is stronger than the conditions we impose in Theorem 2.1 for the Thresholded Lasso estimator. We now show the oracle inequalities for the Dantzig selector. We then show in the supplementary Lemma H.1 that thresholding at the level of $\sigma\lambda$ as elaborated in Step 1 in the Gauss-Dantzig Procedure selects a set $I$ of at most $2s_{0}$ variables, among which at most $s_{0}$ are from the complement of the support set $S$ as required in (155).
Proposition G.4.
@@
Appendix H Proof of Theorem G.3
Now similar to Lemma 2.3, Lemma H.1 bounds the size of $I$, as well as the bias that we introduce to model $I$ by thresholding. Theorem G.3 is an immediate corollary of Lemmas 2.5 and H.1. The proof follows from Lemma H.1 and Proposition G.4. We include its proof in Section H.1 for self-containment.
Lemma H.1.
@@
Proof of Theorem G.3. It holds by definition of $S_{{\mathcal{D}}}$ that $I\cap S_{{\mathcal{D}}}=\emptyset$. It is clear by Lemma H.1 that $|{S_{{\mathcal{D}}}}|<s$ and $|I|\leq 2s_{0}$ and $|I\cup S_{{\mathcal{D}}}|\leq|I\cup S|\leq s+s_{0}\leq 2s$; Thus for $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$, $\hat{\beta}_{I^{c}}=0$ and $\lambda=\sqrt{2\log p/n}$, we have by Lemma 2.7
$\left\lVert\hat{\beta}-\beta\right\rVert_{2}^{2}\leq\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}\left(1+\frac{2\theta^{2}_{s,2s_{0}}}{\Lambda_{\min}^{2}(2s_{0})}\right)+\frac{4s_{0}}{\Lambda_{\min}^{2}(2s_{0})}\lambda_{\sigma,a,p}^{2}$
@@
The proof of Proposition G.4 [9] yields the following on ${\mathcal{T}}_{a}$,
@@
The rest of the proof follows that of Lemma 2.3 and hence omitted. ∎
@@
Figure 5: Illustrative example: i.i.d. Gaussian ensemble; $p=256$, $n=72$, $s=8$, and $\sigma=\sqrt{s}/3$. (a) compare with the Lasso estimator $\tilde{\beta}$ which minimizes $\ell_{2}$ loss. Here $\tilde{\beta}$ has only 3 FPs, but $\rho^{2}$ is large with a value of $64.73$. (b) Compare with the $\beta_{\text{\rm init}}$ obtained using $\lambda_{n}$. The dotted lines show the thresholding level $t_{0}$. The $\beta_{\text{\rm init}}$ has 15 FPs, all of which were cut after the 2nd step; resulting $\rho^{2}=12.73$. After refitting with OLS in the 3rd step, for the $\hat{\beta}$, $\rho^{2}$ is further reduced to $0.51$.
Appendix I Numerical experiments
@@
The experiment set up here is the same as the one in Section  5.1. The main difference is that all non-zero entries in $\beta$ have large magnitudes around 1, in particular, they follow this distribution $\beta_{i}=\mu_{i}(1+|g_{i}|),$ where $\mu_{i}=\pm 1$ with probability 1/2 and $g_{i}\sim N(0,1)$. We use $\lambda_{n}=0.69\lambda\sigma$ throughout the experiments in this section to select a $\beta_{\text{\rm init}}$ as the initial estimator. We then threshold the $\beta_{\text{\rm init}}$ using a value $t_{0}$ typically chosen between $0.5\lambda\sigma$ and $\lambda\sigma$. See each experiment for the actual value used. Given that columns of $X$ being normalized to have $\ell_{2}$ norm $\sqrt{n}$, for each input $\beta$, we compute its SNR as follows: $SNR:=\left\lVert\beta\right\rVert_{2}^{2}/\sigma^{2}.$ To evaluate $\hat{\beta}$, we use metrics defined in Table 3; we also compute the ratio between squared $\ell_{2}$ error and the ideal mean squared error, known as the $\rho^{2}$; see Section I.2 for details.
I.1 Type I/II errors
We now evaluate the Thresholded Lasso estimator by comparing Type I/II errors under different values of $t_{0}$ and SNR. We consider Gaussian random matrices for the design $X$ with both diagonal and Toeplitz covariance. We refer to the former as i.i.d. Gaussian ensemble and the latter as Toeplitz ensemble. In the Toeplitz case, the covariance is given by $T(\gamma)_{i,j}=\gamma^{|i-j|}$ where $0<\gamma<1.$ We run under two noise levels: $\sigma=\sqrt{s}/3$ and $\sigma=\sqrt{s}$. For each $\sigma$, we vary the threshold $t_{0}$ from $0.01\lambda\sigma$ to $1.5\lambda\sigma$. For each $\sigma$ and $t_{0}$ combination, we run the experiment as described in Section 5.1 $200$ times with a new $\beta$ and $\epsilon$ generated each time and we count the number of Type I and II errors in $\hat{\beta}$. We compute the average at the end of 200 runs, which will correspond to one data point on the curves in Figure 6 (a) and (b).
@@ I.1 Type I/II errors
False Positive Rate (FPR) FPR=FP/(FP+TN)=FP/(ps)FPR=FP/(FP+TN)=FP/(p-s)
@@ I.2 $\ell_{2}$ loss
Figure 6: $p=256$ $s=8$. (a) (b) Type I/II errors for i.i.d. Gaussian and Toeplitz ensembles. Each vertical bar represents $\pm 1$ std. The unit of $x$-axis is in $\lambda\sigma$. For both types of design matrices, FPs decrease and FNs increase as the threshold increases. For Toeplitz ensembles, in (c) with fixed correlation $\gamma$, FNs decrease with more samples, and in (d) with fixed sample size, FNs decrease as the correlation $\gamma$ decreases. (e) (f) Histograms of $\rho^{2}$ under i.i.d Gaussian ensembles from 500 runs.
We first run the above experiment using i.i.d. Gaussian ensemble under the following thresholds: $t_{0}=\lambda\sigma$ for $\sigma=\sqrt{s}/3$, and $t_{0}=0.36\lambda\sigma$ for $\sigma=\sqrt{s}$. These are chosen based on the desire to have low errors of both types (as shown in Figure 6 (a)). Naturally, for low SNR cases, small $t_{0}$ will reduce Type II errors. In practice, we suggest using cross-validations to choose the exact constants in front of $\lambda\sigma$; See, for example, a subsequent work [50] for details. We plot the histograms of $\rho^{2}$ in Figure 6 (e) and (f). In (e), the mean and median are $1.45$ and $1.01$ for the Thresholded Lasso, and $46.97$ and $41.12$ for the Lasso. In (f), the corresponding values are $7.26$ and $6.60$ for the Thresholded Lasso and $10.50$ and $10.01$ for the Lasso. With high SNR, the Thresholded Lasso performs extremely well; with low SNR, the improvement of the Thresholded Lasso over the ordinary Lasso is less prominent; this is in close correspondence with the Gauss-Dantzig selector’s behavior as shown by [9].
Next we run the above experiment under different sparsity values of $s$. We again use i.i.d. Gaussian ensemble with $p=2000$, $n=400$, and $\sigma=\sqrt{s}/3$. The threshold is set at $t_{0}=\lambda\sigma$. The SNR for different $s$ is fixed at around $32.36$. Table 4 shows the mean of the $\rho^{2}$ for the Lasso and the Thresholded Lasso estimators. The Thresholded Lasso performs consistently better than the ordinary Lasso until about $s=80$, after which both break down. For the Lasso, we always choose from the full regularization path the optimal $\tilde{\beta}$ that has the minimum $\ell_{2}$ loss.
@@
I.3 Linear Sparsity
We next present results demonstrating that the Thresholded Lasso recovers a sparse model using a small number of samples per non-zero component in $\beta$ when $X$ is a subgaussian ensemble. We run under three cases of $p=256,512,1024$; for each $p$, we increase the sparsity $s$ by roughly equal steps from $s={0.2p}/{\log(0.2p)}$ to $p/4$. For each $p$ and $s$, we run with different sample size $n$. For each tuple $(n,p,s)$, we run an experiment similar to the one described in Section I.1 with an i.i.d. Gaussian ensemble $X$ being fixed while repeating Steps 232-3 100 times. In Step 2, each randomly selected non-zero coordinate of $\beta$ is assigned a value of $\pm 0.9$ with probability $1/2$. After each run, we compare $\hat{\beta}$ with the true $\beta$; if all components match in signs, we count this experiment as a success. At the end of the 100 runs, we compute the percentage of successful runs as the probability of success. We compare with the ordinary Lasso, for which we search over the full regularization path of LARS and choose the $\breve{\beta}$ that best matches $\beta$ in terms of support.
We experiment with $\sigma=1$ and $\sigma=\sqrt{s}/3$. The results are shown in Figure 7. We observe that under both noise levels, the Thresholded Lasso estimator requires much fewer samples than the ordinary lasso in order to conduct exact recovery of the sparsity pattern of the true linear model when all non-zero components are sufficiently large. When $\sigma$ is fixed as $s$ increases, the SNR is increasing; the experimental results illustrate the behavior of sparse recovery when it is close to the noiseless setting. Given the same sparsity, more samples are required for the low SNR case to reach the same level of success rate. Similar behavior was also observed for Toeplitz and Bernoulli ensembles with i.i.d. $\pm 1$ entries.
@@ I.3 Linear Sparsity
Figure 7: (a) (b) Compare the probability of success for $p=256$ and $p=512$ under two noise levels. The Thresholded Lasso estimator requires much fewer samples than the ordinary Lasso. (c) (d) (e) show the probability of success of the Thresholded Lasso under different levels of sparsity and noise levels when $n$ increases for $p=512$ and $1024$. (f) The number of samples $n$ increases almost linearly with $s$ for p = 1024. More samples are required to achieve the same level of success when $\sigma=\sqrt{s}/3$ due to the relatively low SNR.
I.4 ROC comparison
We now compare the performance of the Thresholded Lasso estimator with the Lasso and the Adaptive Lasso by examining their ROC curves. Our parameters are $p=512$, $n=330$, $s=64$ and we run under two cases: $\sigma=\sqrt{s}/3$ and $\sigma=\sqrt{s}$. In the Thresholded Lasso, we vary the threshold level from $0.01\lambda\sigma$ to $1.5\lambda\sigma$. For each threshold, we run the experiment described in Section I.1 with an i.i.d. Gaussian ensemble $X$ being fixed while repeating Steps 232-3 100 times. After each run, we compute the FPR and TPR of the $\hat{\beta}$, and compute their averages after 100 runs as the FPR and TPR for this threshold. For the Lasso, we compute the FPR and TPR for each output vector along its entire regularization path. For the Adaptive Lasso, we use the optimal output $\tilde{\beta}$ in terms of $\ell_{2}$ loss from the initial Lasso penalization path as the input to its second step, that is, we set $\beta_{\text{\rm init}}:=\tilde{\beta}$ and use $w_{j}=1/\beta_{\text{\rm init},j}$ to compute the weights for penalizing those non-zero components in $\beta_{\text{\rm init}}$ in the second step, while all zero components of $\beta_{\text{\rm init}}$ are now removed. We then compute the FPR and TPR for each vector that we obtain from the second step’s LARS output. We implement the algorithms as given in  [52], the details of which are omitted here as its implementation has become standard. The ROC curves are plotted in Figure 8. The Thresholded Lasso performs better than both the ordinary Lasso and the Adaptive Lasso; its advantage is more apparent when the SNR is high.